# Highly Anisotropic Charge Dynamics and Spectral Weight Redistribution in the Trilayer Nickelate La4Ni3O10

### v1

@@ Abstract
We study the abab-plane and $c$-axis charge dynamics of La4Ni3O10 using optical spectroscopy. While a pronounced Drude profile, i.e. metallic response, is observed in the abab-plane optical conductivity $\sigma_{1}^{ab}(\omega)$, the $c$-axis optical spectra $\sigma_{1}^{c}(\omega)$ exhibit semiconducting behavior. The zero-frequency extrapolation of the optical conductivity $\sigma_{1}(\omega\rightarrow 0)\equiv 1/\rho_{\text{dc}}$ gives a resistivity anisotropy of $\rho_{c}/\rho_{ab}\simeq 366$ at 300 K for La4Ni3O10, which is much larger than the values in iron-based superconductors but comparable to those in high-$T_{c}$ cuprates. The interband response is also highly anisotropic, showing salient orbital selectivity for light polarized in the abab plane and along the $c$ axis. The interband-transition peaks in both $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ are located at lower energies compared to density-functional-theory predictions, signifying considerable electronic correlations. By investigating the spectral weight transfer, we find that in the pristine phase, Coulomb correlations have a marked impact on the charge dynamics of La4Ni3O10, whereas in the density-wave state, a gap opens with the Ni-$d_{z^{2}}$ orbital being involved.
@@
Figure 1: (a) Reflectivity of La4Ni3O10 at different temperatures for the abab plane $R^{ab}(\omega)$ and $c$ axis $R^{c}(\omega)$. The inset shows the temperature dependence of $R^{ab}(\omega)$ (red solid circle) and $R^{c}(\omega)$ (blue solid diamond) at $\omega$ = 1000 $\textrm{cm}^{-1}$. (b) The abab-plane optical conductivity $\sigma^{ab}_{1}(\omega)$ and $c$-axis optical conductivity $\sigma^{c}_{1}(\omega)$ of La4Ni3O10 at different temperatures. The inset displays $\sigma^{ab}_{1}(\omega)$ and $\sigma^{c}_{1}(\omega)$ at 300 K on the logarithmic scale with the zero-frequency extrapolations being denoted by solid circles. (c) and (d) show $R(\omega)$ and $\sigma_{1}(\omega)$ at 300 K along different directions in a broad frequency range up to 25 000 $\textrm{cm}^{-1}$, respectively.
@@
We investigate the abab-plane and $c$-axis optical properties of La4Ni3O10. In the abab plane, we found metallic behavior, as evidenced by a conspicuous Drude profile in the abab-plane optical conductivity $\sigma_{1}^{ab}(\omega)$. In sharp contrast, the absence of a Drude response in the $c$-axis optical conductivity $\sigma_{1}^{c}(\omega)$ indicates semiconducting behavior along the $c$ axis. Since $\sigma_{1}(\omega\rightarrow 0)\equiv 1/\rho_{\text{dc}}$, by extrapolating $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ to zero frequency, we obtained $\rho_{c}/\rho_{ab}\simeq 366$ at 300 K. Such a large resistivity anisotropy is reminiscent of high-$T_{c}$ cuprates. Light polarized in the abab plane and along the $c$ axis selectively excites interband transitions between different orbitals, giving rise to a strong anisotropy of the interband response. A detailed examination of the spectral weight redistribution in $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ reveals that the charge dynamics in the pristine phase is strongly influenced by Coulomb correlations, whereas the density-wave state is characterized by the opening of a gap which involves the Ni-$d_{z^{2}}$ orbital.
@@
Sample growth, characterization, optical measurements and data analysis methods are detailed in the Supplemental Materials Sup . Transport measurements indicate that the intertwined charge- and spin-density-wave transition occurs at $T_{\text{DW}}$ = 136.5 K in our La4Ni3O10 sample (Supplemental Materials Sup ). Figure 1(a) shows the reflectivity of La4Ni3O10 for light polarized in the abab plane ($E\parallel ab$) $R^{ab}(\omega)$ and along the $c$ axis ($E\parallel c$) $R^{c}(\omega)$ up to 2000 $\textrm{cm}^{-1}$ at several representative temperatures. $R^{ab}(\omega)$ is high and approaches unity in the low-frequency limit. Additionally, it increases as $T$ is lowered. All these properties suggest a metallic nature of La4Ni3O10 in the abab plane. Below $T_{\text{DW}}$, $R^{ab}(\omega)$ is significantly suppressed in a broad frequency range above 500 $\textrm{cm}^{-1}$ due to the density-wave gap opening, in agreement with a previous optical study Xu et al. (2025). By contrast, $R^{c}(\omega)$ attains a constant value much lower than 1 ($\sim$0.6) in the low-frequency limit and is dominated by a series of sharp features arising from IR-active phonon modes in the far-infrared range, which are prototypical characteristics of semiconducting materials Jiang et al. (2020); Hao et al. (2021); Lobo et al. (2007). The inset of Fig. 1(a) plots the $T$ dependence of $R^{ab}(\omega)$ (red solid circles) and $R^{c}(\omega)$ (blue solid diamonds) at $\omega$ = 1000 $\textrm{cm}^{-1}$. A drop is observed below $T_{\text{DW}}$ in both $R^{ab}(\omega)$ and $R^{c}(\omega)$, indicating that both the abab-plane and $c$-axis charge dynamics are affected by the density-wave transition in La4Ni3O10. Figure 1(c) displays $R^{ab}(\omega)$ (red solid curve) and $R^{c}(\omega)$ (blue solid curve) at 300 K up to 25 000 $\textrm{cm}^{-1}$, highlighting the strong anisotropy of La4Ni3O10 in a broad frequency range.
@@
Figure 1(b) depicts the abab-plane optical conductivity $\sigma_{1}^{ab}(\omega)$ and $c$-axis optical conductivity $\sigma_{1}^{c}(\omega)$ of La4Ni3O10 up to 2000 $\textrm{cm}^{-1}$ at various temperatures. While the low-frequency $\sigma_{1}^{ab}(\omega)$ is characterized by a pronounced Drude response (a peak centered at zero frequency) which is a hallmark of metallic behavior, no Drude response is observed in $\sigma_{1}^{c}(\omega)$. Instead, the far-infrared $\sigma_{1}^{c}(\omega)$ is dominated by sharp phonon peaks, indicating a semiconducting nature of La4Ni3O10 along the $c$ axis. Such semiconducting behavior in $\sigma_{1}^{c}(\omega)$ may be ascribed to the existence of insulating rock salt LaO layers, which effectively prevent the formation of coherence between different trilayers along the $c$ directions. The striking difference between $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ points to highly anisotropic charge dynamics in La4Ni3O10. The zero-frequency extrapolation of $\sigma_{1}(\omega)$ yields the dc conductivity $\sigma_{1}$($\omega\rightarrow$ 0) $\equiv 1/\rho_{\textrm{dc}}$. The inset of Fig. 1(b) plots $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ at 300 K on a logarithmic scale; the red and blue solid circles on the vertical axis denote $\sigma_{1}^{ab}$($\omega\rightarrow$0) $\simeq$ 5490 $\Omega^{-1}\textrm{cm}^{-1}$ and $\sigma_{1}^{c}$($\omega\rightarrow$0) $\simeq$ 15 $\Omega^{-1}\textrm{cm}^{-1}$, respectively, which result in a resistivity anisotropy of $\rho_{c}$$\rho_{ab}$ $\simeq$ 366 at 300 K. This value is significantly larger than those of iron-based superconductors (FeSCs), such as FeSe ($\sim$3–4) Vedeneev et al. (2013), LiFeAs ($\sim$1–3) Song et al. (2010), and the BaFe2As2 family ($\sim$2–6) Tanatar et al. (2009a, b); Nakajima et al. (2018), but comparable to those of high-$T_{c}$ cuprates Komiya et al. (2002); Watanabe et al. (1997). Figure 1(d) compares $\sigma^{ab}_{1}(\omega)$ (red solid curve) and $\sigma^{c}_{1}(\omega)$ (blue solid curve) over a broad frequency range at 300 K, revealing that the interband electronic transitions are also strongly anisotropic in La4Ni3O10.
@@
In summary, the abab-plane and $c$-axis charge dynamics of La4Ni3O10 have been studied using optical spectroscopy. While $\sigma_{1}^{ab}(\omega)$ exhibits a pronounced Drude peak, indicating a metallic nature of La4Ni3O10 in the abab plane, semiconducting behavior is observed in $\sigma_{1}^{c}(\omega)$. The zero-frequency extrapolation of $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ at 300 K yields a resistivity anisotropy of $\rho_{c}/\rho_{ab}\simeq 366$ for La4Ni3O10, which is much higher than the values in FeSCs, but similar to those in cuprates. The interband response also features strong anisotropy, as light polarized in the abab plane and along the $c$ axis selectively excites electronic transitions between different orbitals. In the pristine phase, the spectral weight transfer implies that the charge dynamics is under the influence of Coulomb correlations, whereas in the density-wave state, the spectral weight redistribution suggests the opening of a gap with the participation of the Ni-$d_{z^{2}}$ orbital.

### v2

@@ Abstract
We study the $ab$-plane and $c$-axis charge dynamics of La4Ni3O10 using optical spectroscopy. While a pronounced Drude profile, i.e. metallic response, is observed in the $ab$-plane optical conductivity $\sigma_{1}^{ab}(\omega)$, the $c$-axis optical spectra $\sigma_{1}^{c}(\omega)$ exhibit semiconducting behavior. The zero-frequency extrapolation of the optical conductivity $\sigma_{1}(\omega\rightarrow 0)\equiv 1/\rho_{\text{dc}}$ gives a resistivity anisotropy of $\rho_{c}/\rho_{ab}\simeq 366$ at 300 K for La4Ni3O10, which is much larger than the values in iron-based superconductors but comparable to those in high-$T_{c}$ cuprates. The interband response is also highly anisotropic, showing salient orbital selectivity for light polarized in the $ab$ plane and along the $c$ axis. The interband-transition peaks in both $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ are located at lower energies compared to density-functional-theory predictions, signifying considerable electronic correlations. By investigating the spectral weight transfer, we find that in the pristine phase, Coulomb correlations have a marked impact on the charge dynamics of La4Ni3O10, whereas in the density-wave state, a gap opens with the Ni-$d_{z^{2}}$ orbital being involved.
@@
Figure 1: (a) Reflectivity of La4Ni3O10 at different temperatures for the $ab$ plane $R^{ab}(\omega)$ and $c$ axis $R^{c}(\omega)$. The inset shows the temperature dependence of $R^{ab}(\omega)$ (red solid circle) and $R^{c}(\omega)$ (blue solid diamond) at $\omega$ = 1000 $\textrm{cm}^{-1}$. (b) The abab-plane optical conductivity $\sigma^{ab}_{1}(\omega)$ and $c$-axis optical conductivity $\sigma^{c}_{1}(\omega)$ of La4Ni3O10 at different temperatures. The inset displays $\sigma^{ab}_{1}(\omega)$ and $\sigma^{c}_{1}(\omega)$ at 300 K on the logarithmic scale with the zero-frequency extrapolations being denoted by solid circles. (c) and (d) show $R(\omega)$ and $\sigma_{1}(\omega)$ at 300 K along different directions in a broad frequency range up to 25 000 $\textrm{cm}^{-1}$, respectively.
@@
We investigate the $ab$-plane and $c$-axis optical properties of La4Ni3O10. In the $ab$ plane, we found metallic behavior, as evidenced by a conspicuous Drude profile in the $ab$-plane optical conductivity $\sigma_{1}^{ab}(\omega)$. In sharp contrast, the absence of a Drude response in the $c$-axis optical conductivity $\sigma_{1}^{c}(\omega)$ indicates semiconducting behavior along the $c$ axis. Since $\sigma_{1}(\omega\rightarrow 0)\equiv 1/\rho_{\text{dc}}$, by extrapolating $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ to zero frequency, we obtained $\rho_{c}/\rho_{ab}\simeq 366$ at 300 K. Such a large resistivity anisotropy is reminiscent of high-$T_{c}$ cuprates. Light polarized in the $ab$ plane and along the $c$ axis selectively excites interband transitions between different orbitals, giving rise to a strong anisotropy of the interband response. A detailed examination of the spectral weight redistribution in $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ reveals that the charge dynamics in the pristine phase is strongly influenced by Coulomb correlations, whereas the density-wave state is characterized by the opening of a gap which involves the Ni-$d_{z^{2}}$ orbital.
@@
Sample growth, characterization, optical measurements and data analysis methods are detailed in the Supplemental Materials Sup . Transport measurements indicate that the intertwined charge- and spin-density-wave transition occurs at $T_{\text{DW}}$ = 136.5 K in our La4Ni3O10 sample (Supplemental Materials Sup ). Figure 1(a) shows the reflectivity of La4Ni3O10 for light polarized in the $ab$ plane ($E\parallel ab$) $R^{ab}(\omega)$ and along the $c$ axis ($E\parallel c$) $R^{c}(\omega)$ up to 2000 $\textrm{cm}^{-1}$ at several representative temperatures. $R^{ab}(\omega)$ is high and approaches unity in the low-frequency limit. Additionally, it increases as $T$ is lowered. All these properties suggest a metallic nature of La4Ni3O10 in the $ab$ plane. Below $T_{\text{DW}}$, $R^{ab}(\omega)$ is significantly suppressed in a broad frequency range above 500 $\textrm{cm}^{-1}$ due to the density-wave gap opening, in agreement with a previous optical study Xu et al. (2025). By contrast, $R^{c}(\omega)$ attains a constant value much lower than 1 ($\sim$0.6) in the low-frequency limit and is dominated by a series of sharp features arising from IR-active phonon modes in the far-infrared range, which are prototypical characteristics of semiconducting materials Jiang et al. (2020); Hao et al. (2021); Lobo et al. (2007). The inset of Fig. 1(a) plots the $T$ dependence of $R^{ab}(\omega)$ (red solid circles) and $R^{c}(\omega)$ (blue solid diamonds) at $\omega$ = 1000 $\textrm{cm}^{-1}$. A drop is observed below $T_{\text{DW}}$ in both $R^{ab}(\omega)$ and $R^{c}(\omega)$, indicating that both the $ab$-plane and $c$-axis charge dynamics are affected by the density-wave transition in La4Ni3O10. Figure 1(c) displays $R^{ab}(\omega)$ (red solid curve) and $R^{c}(\omega)$ (blue solid curve) at 300 K up to 25 000 $\textrm{cm}^{-1}$, highlighting the strong anisotropy of La4Ni3O10 in a broad frequency range.
@@
Figure 1(b) depicts the $ab$-plane optical conductivity $\sigma_{1}^{ab}(\omega)$ and $c$-axis optical conductivity $\sigma_{1}^{c}(\omega)$ of La4Ni3O10 up to 2000 $\textrm{cm}^{-1}$ at various temperatures. While the low-frequency $\sigma_{1}^{ab}(\omega)$ is characterized by a pronounced Drude response (a peak centered at zero frequency) which is a hallmark of metallic behavior, no Drude response is observed in $\sigma_{1}^{c}(\omega)$. Instead, the far-infrared $\sigma_{1}^{c}(\omega)$ is dominated by sharp phonon peaks, indicating a semiconducting nature of La4Ni3O10 along the $c$ axis. Such semiconducting behavior in $\sigma_{1}^{c}(\omega)$ may be ascribed to the existence of insulating rock salt LaO layers, which effectively prevent the formation of coherence between different trilayers along the $c$ directions. The striking difference between $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ points to highly anisotropic charge dynamics in La4Ni3O10. The zero-frequency extrapolation of $\sigma_{1}(\omega)$ yields the dc conductivity $\sigma_{1}$($\omega\rightarrow$ 0) $\equiv 1/\rho_{\textrm{dc}}$. The inset of Fig. 1(b) plots $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ at 300 K on a logarithmic scale; the red and blue solid circles on the vertical axis denote $\sigma_{1}^{ab}$($\omega\rightarrow$0) $\simeq$ 5490 $\Omega^{-1}\textrm{cm}^{-1}$ and $\sigma_{1}^{c}$($\omega\rightarrow$0) $\simeq$ 15 $\Omega^{-1}\textrm{cm}^{-1}$, respectively, which result in a resistivity anisotropy of $\rho_{c}$$\rho_{ab}$ $\simeq$ 366 at 300 K. This value is significantly larger than those of iron-based superconductors (FeSCs), such as FeSe ($\sim$3–4) Vedeneev et al. (2013), LiFeAs ($\sim$1–3) Song et al. (2010), and the BaFe2As2 family ($\sim$2–6) Tanatar et al. (2009a, b); Nakajima et al. (2018), but comparable to those of high-$T_{c}$ cuprates Komiya et al. (2002); Watanabe et al. (1997). Figure 1(d) compares $\sigma^{ab}_{1}(\omega)$ (red solid curve) and $\sigma^{c}_{1}(\omega)$ (blue solid curve) over a broad frequency range at 300 K, revealing that the interband electronic transitions are also strongly anisotropic in La4Ni3O10.
@@
In summary, the $ab$-plane and $c$-axis charge dynamics of La4Ni3O10 have been studied using optical spectroscopy. While $\sigma_{1}^{ab}(\omega)$ exhibits a pronounced Drude peak, indicating a metallic nature of La4Ni3O10 in the $ab$ plane, semiconducting behavior is observed in $\sigma_{1}^{c}(\omega)$. The zero-frequency extrapolation of $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ at 300 K yields a resistivity anisotropy of $\rho_{c}/\rho_{ab}\simeq 366$ for La4Ni3O10, which is much higher than the values in FeSCs, but similar to those in cuprates. The interband response also features strong anisotropy, as light polarized in the $ab$ plane and along the $c$ axis selectively excites electronic transitions between different orbitals. In the pristine phase, the spectral weight transfer implies that the charge dynamics is under the influence of Coulomb correlations, whereas in the density-wave state, the spectral weight redistribution suggests the opening of a gap with the participation of the Ni-$d_{z^{2}}$ orbital.